# Modeling Extensive Defects in Metals through Active Machine Learning and Automated Configuration Reconstruction

### v1

@@ DISCUSSION
While the MLIP-3 package facilitates OTF-AL using MTPs to tackle similar challenges  (?), our integration of EIP-AL and PCC-GCMC offers substantial improvements. EIP-AL significantly reduces the demanding requirements of OTF-AL, expediting the development of MLIPs. Typically, OTF-AL involves iterative MD simulations based on MLIPs and frequent retraining  (?), a process that is time-consuming, especially for rare events like diffusion and phase transitions. Moreover, retraining MLIPs—whether through linear regression or neural networks—is costly and computationally intensive. The data in Table S1 highlights this efficiency. Of the 348 configurations generated, only 72 are from OTF-AL. On the other hand, the new LAEDs acquired from OTF-AL are usually specific to the simulated loading conditions and may not generalize well across different scenarios. It is impractical to perform OTF-AL for every conceivable loading condition. For example, certain LAEDs that occur in low strain-rate MD simulations might not emerge under high strain rates, such as the rate-dependent ductile-to-brittle failure transition observed in metal nanowires  (?, ?). Consequently, OTF-AL from MD simulations at high strain rates cannot be used to develop MLIPs suitable for lower strain rates, as critical LAEDs at lower rates remain unexplored, as evidenced in. This strain rate limitation has been identified as a significant challenge in advancing MLIP development  (?). EIP-AL effectively addresses this challenge by leveraging cost-effective EIPs to conduct extensive MD simulations. Even within the same material, various EIPs exhibit distinct properties, which can produce a diverse range of LAEDs under identical loading conditions, as demonstrated by the crack propagation scenarios shown in Fig. 4. For tungsten alone, there are 32 different EIPs available from OpenKIM. The varied characteristics of these EIPs allow us to discover a more extensive array of LAEDs than those typically identified through OTF-AL.
@@
Our PCC-GCMC method significantly refines our approach to developing MLIPs for modeling extensive defects, serving as a critical supplement to both EIP-AL and OTF-AL. This method departs from the conventional approach used in the MLIP-3 package  (?), which involves adding a vacuum layer to minimize image interactions—a process that proves inefficient due to the limited use of LAEs, with most being near the surface (Fig. 2a). Additionally, this traditional approach requires further DFT calculations for each LAED using the Gamma point, a necessity due to the increased box size following the introduction of the vacuum layer. This complicates the training process because of the resulting inconsistent Kpoint density (Fig. 2b). In contrast, PCC-GCMC streamlines the creation of periodic configurations that are immediately integrated into the training dataset following DFT calculations, enhancing the dataset’s relevance, as discussed in Fig. 2 and Fig. S2. Specifically, This method avoids the use of a vacuum layer, circumventing issues with surface atoms. By selecting only the lowest-energy configurations from PCC-GCMC, we ensure that all LAEs in these configurations are physically pertinent to the central LAED. The introduction of atoms in this process could yield perfect lattice or related defect structures (Fig. S1), or more likely, identify previously unrecognized defects that improve the transferability of the developed MLIP (Fig. S2).
@@
Finally, we propose that our framework can be generalized across a diverse array of common crystalline materials to model extensive defects, as illustrated in Fig. 7. This framework is versatile, applicable not only to familiar structures such as FCC, BCC, HCP, and diamond cubic but also to more complex lattices like perovskite. A rich assortment of EIPs is available from resources like the NIST Interatomic Potentials Repository (NIST-IPR)  (?) and OpenKIM  (?). These resources provide access to a variety of potentials, including Embedded Atom Model (EAM), Modified Embedded Atom Method (MEAM), bond order potentials (BOP), Stillinger-Weber (SW) potentials, Tersoff potentials, and reactive force fields (ReaxFF), among others. OpenKIM, for instance, hosts over 644 potentials, with a significant majority being EIPs. Computational materials science has benefited greatly over the past three decades by leveraging these EIPs  (?), and despite the current emphasis on MLIPs, the value of EIPs remains substantial. Although these potentials exhibit higher force and energy errors compared to DFT calculations, their consistent internal logic as physical models provides significant value. Utilizing the abundant EIPs available, our study’s approach can be used to develops advanced MLIPs through three strategic initiatives: adapting EIPs from different materials but sharing the same lattice structure by merely rescaling lattice constants to generate relevant LAEDs; parameterizing EIPs to customize the mechanical properties of materials  (?, ?); and employing the average-atom interatomic potential for random alloys to produce an unlimited series of EIPs from basic elemental forms  (?). Utilizing these resources, our approach enables the creation of a standard defect genome for each crystalline structure. These genomes form the cornerstone for developing MLIPs capable of accurately modeling any type of defect. This not only facilitates further research within this study but also provides a valuable tool for the broader scientific community, enhancing the predictive capabilities of computational material science.
Figure 7: Implications of this study on MLIP development for modeling extensive defects in general crystalline materials.
@@
In summary, this work introduces an advanced framework for developing MLIPs by integrating EIP-AL, PCC-GCMC, and OTF-AL. This framework enables automated, ab initio-accurate, large-scale atomistic modeling of general defects in metals. The developed MLIPs are highly transferable and are based on comprehensive datasets that facilitate simulations of complex behaviors, including general GBs in polycrystals, dislocation nucleation and interactions, dislocation-GB interactions, and crack propagation in BCC W. Rigorous uncertainty quantification analyses, employing D-optimality criterion and ensemble learning techniques, confirm the accuracy of our models in these scenarios. We have successfully applied our trained ACE potential to uncover dominant deformation mechanisms in nanoindentation, identifying deformation twinning as the primary mechanism at the onset of plasticity, and dislocation evolution as the predominant mechanism at later stages. Additionally, we have elucidated the distinct strengthening mechanisms between low-angle and high-angle twist GBs in spherical bicrystal compression tests. This approach proves versatile enough to be generalized for modeling extensive defects in other BCC metals, FCC metals, and high-entropy alloys, demonstrating its broad applicability across various material systems.
@@
We employ the MLIP-2 package  (?) and pacemarker  (?) to develop MTPs and ACE potential for BCC W. MLIP-2 utilizes moment tensor descriptors and apply linear regression to train the machine learning model, enabling it to predict the energy, force, and stress of atomistic systems. Within our study, the weights of energy, force and stress are set as 1, 0.01 and 0, respectively. For ACE, we utilize the highly nonlinear potential energy function with the complex form of $E_{i}=\varphi+\sqrt{\varphi}+\sum_{i}\varphi^{f_{i}}$, where $E_{i}=\varphi+\sqrt{\varphi}+\sum_{i}\varphi^{f_{i}},\quad f_{i}\in\left\{\frac{1}{8},\frac{1}{4},\frac{3}{8},\frac{3}{4},\frac{7}{8},2\right\}$, obtained from Ref  (?). For the expansion of the atomic properties, we used 72 basis functions with 801 parameters. As radial basis we employ Bessel functions. A $\kappa$ value of 0.01, which gives the ratio between force and energy weights value, was used during fitting. For optimisation we use the BFGS algorithm for 2000 steps. The cutoff distance is set as 5 Å for both MTP and ACE. More importantly, we utilize MLIP-3  (?) to perform active learning based on the MaxVol algorithm from extensive simulations conducted with EAM or MEAM potentials. This method was originally developed for the OFT-AL of atomic environments, incorporating a cycle of MTP-based simulations, DFT calculations, and subsequent MTP retraining. In this study, to expedite the process, we exclusively utilize the EIP-AL to identify representative defects from a comprehensive pool of large-scale simulations. This approach and its implications are further elaborated in Section 3.
@@
The data that support the findings of this study are available upon reasonable request. The parameter files of the potential and training dataset will be published upon the publication of this paper.
@@
All simulations are executed using open-source software LAMMPS. The machine learning force field was trained and validated by the MLIP package (?).
@@ Acknowledgments
Modeling Extensive Defects in Metals through Active Machine Learning and Automated Configuration Reconstruction
@@
Figure S1 to S6
Supplementary Note 1. PCC-GCMC for a screw dislocation core
Fig. S1 utilizes the example of a screw dislocation core in BCC W to demonstrate the PCC-GCMC process. As shown in Fig. S1a, the atomic cluster includes a screw dislocation core with three compact core atoms, labeled 1-3. This configuration maintains periodic boundary conditions along the z-direction, with free boundary conditions along the $x$ and $y$ directions. The PCC-GCMC process is repeated with varying box dimensions, $\Delta_{x}$ and $\Delta_{y}$, from 1 to 10 Å. Fig. S1b displays the minimum potential energies per atom post-PCC-GCMC, indicating that the resulting periodic configurations exhibit varied potential energies. Notably, configurations with potential energies below -8.8 eV/atom, depicted in blue, approach the cohesive energy of -8.9 eV/atom and are nearly equivalent to a perfect BCC lattice. Fig. S1c highlights two representative configurations. In the top panel, we observe a periodic configuration featuring a new non-compact core crossing the periodic boundary along the $y$-axis dislocation, with the remainder of the atoms conforming to the perfect BCC lattice structure, excluding the original screw core. The bottom panel reveals a periodic configuration that includes a screw dislocation dipole, akin to typical screw dislocation dipole configurations with a tilted box. Both configurations depicted in Fig. S1c are generated entirely through automated processes. All newly inserted atoms either form a pure BCC lattice or constitute relevant LAEs, making them suitable for direct inclusion in our MLIP training.
@@
Fig. S4 showcases the uncertainty quantification results of a polycrystal relaxed at 300 K and 2000 K, based on the collective datasets of B+P and on-the-fly AL for GBs (OTF-AL-GB) sessions to train five independent MTPs. We then calculate the standard deviation of atomic forces ($\sigma_{F}$) for all atoms according to Equation S1. Comparative calculations are also performed using five independent MTPs trained using only the basic dataset (B). Figs S4a, c show the GB atoms color-coded by $\sigma_{F}$ at 300 K and 2000 K based on dataset B and the combined datasets B+P+OTF-AL-GB, respectively. Notably, a significant number of GB atoms exhibit $\sigma_{F}$ values higher than 0.5 eV/Å at both temperatures when relying solely on dataset B. In contrast, all atoms show $\sigma_{F}$ values below 0.5 eV/Å at both the temperatures when assessed with the combined datasets. The histogram plots in Figs S4b, d further validate that the MLIPs trained on the collective datasets (B+P+OTF-AL-GB) can accurately predict the atomic forces for all GBs at both 300 K and 2000 K, highlighting the enhanced precision and reliability of our model in high-temperature conditions.
Supplementary Note 3. Final MLIP training and selection
@@
We possess a comprehensive final database that aggregates representative LAEDs in periodic configurations: B+P+OTF-AL-GB+EIP-AL-fracture+OTF-AL-fracture. Detailed breakdowns of each dataset are presented in Table S1. Altogether, the database encompasses 1026 configurations and 94,516 atoms. Of the 348 configurations newly incorporated, 276 are obtained through EIP-AL, while only 72 results from on-the-fly AL. This distribution highlights the efficiency of EIP-AL in significantly reducing the reliance on the more resource-intensive OTF-AL, thereby expediting the development of MLIPs.
@@
The MLIPs trained from this database are equipped to simulate nearly all plastic deformations and fracture of BCC W. However, configurations necessary for radiation simulations due to cascade displacement are currently not included. These configurations could easily be integrated from existing databases or generated through EIP-AL and OTF-AL in future enhancements. We utilize different levels of MTP and a complex ACE potential for training various MLIPs. We particularly favor ACE due to its computational efficiency and the ease of calculating extrapolation grades during simulations in LAMMPS, which greatly facilitates the ongoing assessment. The analysis of basic properties of BCC W, computational efficiency, and training/test performance are detailed in Table S2. It is observed that the accuracy of MTP models increases with their complexity level, but at the cost of reduced computational efficiency. Specifically, the computational expense of MTP20 is seven times that of MTP12. In contrast, ACE demonstrates accuracy comparable to MTP20 while maintaining efficiency similar to MTP12. Based on these findings, ACE is selected as the optimal model for modeling general defects in BCC W due to its superior balance of accuracy and efficiency.
@@
$N_{\text{cfg}}$ Average atom number
Basic 678 72
Initial EIP-AL 262 133
OTF-AL-GB 47 135
EIP-AL-fracture 14 118
OTF-AL-fracture 25 115
Total 1026 92
@@
Figure S1: Illustration of PCC-GCMC using a screw dislocation core. a Atomic cluster showcasing a screw dislocation core; atoms are color-coded based on Common Neighbor Analysis (CNA) shown in the bottom panel. b Average potential energy per atom for each configuration post atom insertion. c Two representative periodic configurations corresponding to points A and B in b.
Figure S2: Direct validation of MLIPs in nice grain boundaries (GBs) by comparing energy and force from DFT calculations. a Pristine GB structures. b Compressed GB structures. All atoms in a and b are color-coded by common neighbor analysis. c Energy comparison. d Atomic force comparison.
@@
Figure S4: Uncertainty quantification based on ensemble learning for a polycrystal. a Results from the basic dataset. b The probability density function of $\sigma_{F}$ at temperatures of 300 K and 2000 K, derived from the basic dataset. c Results from the combined dataset (B+P+OTF-GB). d The probability density function of $\sigma_{F}$ at 300 K and 2000 K based on the combined dataset.
Figure S5: Additional plots of two crystals compression. a The evolution of contact area with respect to displacement for two bicrystals. b The evolution of force with respect to displacement for two bicrystals. c The evolution of the maximum extrapolation grade ($\gamma_{\text{max}}$) with respect to displacement for two bicrystals.
Figure S6: Extracted dislocation lines during bicrystal compression simulations. A-D correspond to the critical points in Fig. 6c. Left panels are snapshots of bicrystal containing a low-angle twist GB, and right panels are snapshots of the bicrystal containing a high-angle twist GB.

### v2

@@ DISCUSSION
While the MLIP-3 package facilitates OTF-AL using MTPs to tackle similar challenges  (?), our integration of EIP-GS and PCC-GCMC offers substantial improvements. EIP-GS significantly reduces the demanding requirements of OTF-AL, expediting the development of MLIPs. Typically, OTF-AL involves iterative MD simulations based on MLIPs and frequent retraining  (?), a process that is time-consuming, especially for rare events like diffusion and phase transitions. Moreover, retraining MLIPs, whether using linear regression or neural networks, is computationally intensive. The data in Table S1 highlights this efficiency. Of the 348 configurations generated, only 72 are from OTF-AL. On the other hand, the new LAEs acquired from OTF-AL are usually specific to the simulated loading conditions and may not generalize well across different scenarios. It is impractical to perform OTF-AL for every conceivable loading condition. For example, certain LAEs that occur in low strain-rate MD simulations might not emerge under high strain rates, such as the rate-dependent ductile-to-brittle failure transition observed in metal nanowires  (?, ?). Consequently, OTF-AL from MD simulations at high strain rates cannot be used to develop MLIPs suitable for lower strain rates, as critical LAEs at lower rates remain unexplored. This strain rate limitation has been identified as a significant challenge in advancing MLIP development  (?). EIP-GS effectively addresses this challenge by leveraging cost-effective EIPs to conduct extensive MD simulations. Even within the same material, various EIPs exhibit distinct properties, which can produce a diverse range of LAEs under identical loading conditions, as demonstrated by the crack propagation scenarios shown in Fig. S2. For W alone, there are 32 different EIPs available from OpenKIM  (?). The varied characteristics of these EIPs allow us to discover a more extensive array of LAEs than those typically identified through OTF-AL.
@@
Our PCC-GCMC method significantly refines our approach to developing MLIPs for modeling extensive defects, serving as a critical supplement to both EIP-GS and OTF-AL. This method departs from the conventional approach used in the MLIP-3 package  (?), which involves adding a vacuum layer to minimize image interactions—a process that proves inefficient due to the limited use of LAEs, with most being near the surface (Fig. 2a). In contrast, PCC-GCMC streamlines the creation of periodic configurations that are immediately integrated into the training dataset following DFT calculations, enhancing the dataset’s relevance, as discussed in Fig. 2. Specifically, this method avoids the use of a vacuum layer, circumventing issues with surface atoms. The introduction of atoms in this process could yield perfect lattice or related defect structures (Fig. 2a), or more likely, identify previously unrecognized defects that improve the transferability of the developed MLIP (Fig. 4).
Our approach demonstrates significant advancements over previous studies  (?, ?, ?). While these earlier methods are restricted to simple defects and often rely on manual operations, such as rotation and merging in Ref.  (?), they also require elastic solutions for simple defects to construct atomistic positions, as seen in Refs.  (?, ?). In contrast, our methodology is uniquely capable of addressing complex defects, including mixed dislocations and general GBs, where prior approaches prove inadequate. An alternative strategy for developing general-purpose MLIP datasets (RANDSPG) involves the systematic exploration of all possible space groups in random crystal structures  (?, ?). Although MLIPs trained on the RANDSPG dataset achieve accurate modeling of simple GBs in Mg and general GBs in BCC Fe  (?), recent findings reveal substantial limitations. Specifically, high uncertainty ($\gamma>2$) and significant force errors are observed in polycrystal deformations of BCC Fe  (?), as reported by the same authors in Ref.  (?). In comparison, our MLIPs exhibit markedly lower uncertainty ($\gamma<1.2$) in nanoindentation (Fig. 6c). These results underscore the insufficiency of the RANDSPG dataset for modeling extensive defects and highlight the superior robustness of our approach.
@@
Finally, we propose that our framework can be generalized across a diverse array of common crystalline materials to model extensive defects, as illustrated in Fig. 7. This framework is versatile, applicable not only to familiar structures such as FCC, BCC, HCP, and diamond cubic but also to more complex lattices like perovskite. A rich assortment of EIPs is available from resources like the NIST Interatomic Potentials Repository (NIST-IPR)  (?) and OpenKIM  (?). These resources provide access to a variety of potentials, including EAM, MEAM, bond order potentials (BOP), Stillinger-Weber (SW) potentials, Tersoff potentials, and reactive force fields (ReaxFF), among others. OpenKIM, for instance, hosts over 644 potentials, with a significant majority being EIPs. Computational materials science has benefited greatly over the past three decades by leveraging these EIPs  (?), and despite the current emphasis on MLIPs, the value of EIPs remains substantial. Although these potentials exhibit higher force and energy errors compared to DFT calculations, their consistent internal logic as physical models provides significant value. Utilizing the abundant EIPs available, our study’s approach can be used to develops advanced MLIPs through three strategic initiatives: adapting EIPs from different materials but sharing the same lattice structure by merely rescaling lattice constants to generate relevant LAEs; parameterizing EIPs to customize the mechanical properties of materials  (?, ?); and employing the average-atom interatomic potential for random alloys to produce an unlimited series of EIPs from basic elemental forms  (?). Utilizing these resources, our approach enables the creation of a standard defect genome for each crystalline structure. These genomes form the cornerstone for developing MLIPs capable of accurately modeling any type of defect. This not only facilitates further research within this study but also provides a valuable tool for the broader scientific community, enhancing the predictive capabilities of computational material science.
@@
In summary, this work introduces an advanced framework for developing MLIPs by integrating EIP-GS, PCC-GCMC, and OTF-AL. This framework enables automated, ab initio-accurate, large-scale atomistic modeling of general defects in metals. The developed MLIPs are highly transferable and are based on comprehensive datasets that facilitate simulations of complex behaviors, including general GBs in polycrystals, dislocation nucleation and interactions, dislocation-GB interactions, and crack propagation in BCC W. Rigorous uncertainty quantification analyses, employing D-optimality criterion and ensemble learning techniques, confirm the accuracy of our models in these scenarios. We have successfully applied our trained ACE potential to uncover dominant deformation mechanisms in nanoindentation, identifying deformation twinning as the primary mechanism at the onset of plasticity, and dislocation evolution as the predominant mechanism at later stages. This approach proves versatile enough to be generalized for modeling extensive defects in other BCC metals, FCC metals, and high-entropy alloys, demonstrating its broad applicability across various material systems.
@@
The indentation is conducted along the [100] direction, which is free from boundary constraints, while the other two axes maintain periodic boundary conditions. The total system spans dimensions of 38 $\times$ 38 $\times$ 32 $\text{nm}^{3}$, encompassing 2,832,200 atoms. To stabilize the substrate, the bottom layers, with a thickness of 6.4 Å, are fixed. Additionally, several layers, totaling a thickness of 2.56 nm, function as thermostat layers to dissipate heat generated during the nanoindentation process. The remaining atoms are designated as Newtonian layers, where the indentation process occurs. The indentation is performed at a velocity of 20 m/s, using a spherical indenter with a radius of 6 nm.
@@
We employ the MLIP-2 package  (?) and pacemarker  (?) to develop MTPs and ACE potential for BCC W. MLIP-2 utilizes moment tensor descriptors and apply linear regression to train the machine learning model, enabling it to predict the energy, force, and stress of atomistic systems. Within our study, the weights of energy, force and stress are set as 1, 0.01 and 0, respectively. For ACE, we utilize the highly nonlinear potential energy function with the complex form of $E_{i}=\varphi+\sqrt{\varphi}+\sum_{i}\varphi^{f_{i}}$, where $E_{i}=\varphi+\sqrt{\varphi}+\sum_{i}\varphi^{f_{i}},\quad f_{i}\in\left\{\frac{1}{8},\frac{1}{4},\frac{3}{8},\frac{3}{4},\frac{7}{8},2\right\}$, obtained from Ref  (?). For the expansion of the atomic properties, we used 72 basis functions with 801 parameters. As radial basis we employ Bessel functions. A $\kappa$ value of 0.01, which gives the ratio between force and energy weights value, was used during fitting. For optimisation we use the BFGS algorithm for 2000 steps. The cutoff distance is set as 5 Å for both MTP and ACE. More importantly, we utilize MLIP-3  (?) to perform active learning based on the MaxVol algorithm from extensive simulations conducted with EAM or MEAM potentials. This method was originally developed for the OFT-AL of atomic environments, incorporating a cycle of MTP-based simulations, DFT calculations, and subsequent MTP retraining. In this study, to expedite the process, we exclusively utilize the EIP-GS to identify representative defects from a comprehensive pool of large-scale simulations. This approach and its implications are further elaborated in Section 3.
@@
The data that support the findings of this study are available at the GitHub repository: https://github.com/ufsf/ML-Defect-Modeling.
@@
All simulations are executed using open-source software LAMMPS. The machine learning force field was trained and validated by the MLIP package (?). All source codes are available at the GitHub repository: https://github.com/ufsf/ML-Defect-Modeling.
@@ Acknowledgments
Modeling Extensive Defects in Metals through Classical Potential-Guided Sampling and Automated Configuration Reconstruction
@@
Figure S1 to S2
Supplementary Note 1. LAEs comparison among datasets B+C, B+P and GAP
The basic dataset contains 678 configurations, which are obtained by domain knowledge. This collection encompasses ground-state structures, configurations deformed under various elastic strains, ab initio molecular dynamics (AIMD) simulations across multiple temperatures, edge and screw dislocations, and simple grain boundaries  (?). These configurations have been curated from our recent work  (?). The GAP dataset is obtained from Ref.  (?), totaling 9693 configurations. The dataset includes primitive unit cell with varying lattice vectors, vacancy, dislocation quadrupole, unstable stacking fault, surface.
To further highlight the greater diversity of LAE in the B+P dataset compared to B+C, we calculate the extrapolation grade $\gamma$ for configurations within the GAP dataset, evaluated by both B+C and B+P. The comparative results are illustrated in Fig. S1a, where we plot the $\gamma$ values for B+C and B+P against each other. Our analysis reveals that the B+P dataset includes 59 configurations with $\gamma$ values exceeding 5, notably, none exceeding 20. In contrast, the B+C dataset shows 74 configurations with $\gamma$ values over 5, including 8 configurations where $\gamma$ surpasses 20. This stark contrast underscores that B+P offers a richer diversity of LAEs than B+C. Interestingly, Fig. S1a also indicates that certain LAEs present in the GAP dataset are missing in B+P. These absent LAEs are primarily configurations of dumbbell interstitials, which are particularly relevant to simulations of radiation damage. Nonetheless, incorporating these specific LAEs into our model through EIP-GS and PCC-GCMC would be straightforward. Furthermore, we calculate the extrapolation grade $\gamma$ for configurations within the dataset B+P based on the dataset GAP, as depicted in Fig. S1b. It is observed that 14.8% of configurations within B+P are absent in GAP. This significant absence explains the high uncertainty observed for the dataset GAP in Fig. 3, indicating that the existing GAP dataset and the resultant MLIP are not adequately equipped for general defect modeling of BCC W. This finding underscores the necessity of our study and the development of a more comprehensive dataset such as B+P, which better addresses the complexities of large-scale simulations.
@@
Supplementary Note 3. MLIP refinement for crack propagation
In this work, we examine crack propagation in BCC W across four distinct crack systems. Our initial EIP-GS model, referenced in Fig. 1a, did not include fracture simulations. We therefore conducted simulations using both EAM-Zhou  (?) and MEAM potentials  (?), widely recognized in fracture analysis. We present key snapshots of crack propagation in Figs S2a-d using EAM and Figs S2e-h using MEAM. MEAM, optimized for superior fracture simulation, exhibits distinct fracture behaviors compared to EAM. For instance, EAM shows a BCC-to-FCC phase transition at the crack tip within the (010)[001] system (Fig. S2a), while MEAM primarily displays brittle cleavage (Fig. S2e), although a similar phase transition is observed in later stages, potentially influenced by boundary conditions. Additionally, EAM captures dislocation emission in the systems $(\bar{1}10)[001]$ and $(\bar{1}10)[\bar{1}\bar{1}1]$ (Figs S2b, d), unlike the brittle cleavage observed with MEAM in these systems (Figs S2f, h). Both potentials, however, show deformation twinning at the crack tip in the system $(112)[1\bar{1}0]$ (Figs S2c, g). To refine our MLIPs, EIP-GS is applied to simulation outputs from both EAM and MEAM, resulting in the selection of 14 new defect clusters essential for fracture simulations. These clusters are processed using PCC-GCMC to create periodic configurations for further DFT calculations, forming the new dataset, EIP-GS-fracture. Building on this, a new MTP is trained incorporating the datasets B+P+OTF-AL-GB+EIP-GS-fracture. Subsequent OTF-AL for fracture simulations across the four systems leads to the selection of 25 additional configurations. After converting these into periodic configurations via PCC-GCMC and processing them with DFT, they are integrated into the comprehensive datasets, culminating in B+P+OTF-AL-GB+EIP-GS-fracture+OTF-AL-fracture.
To demonstrate the reduction in uncertainty achieved by integrating the EIP-GS-fracture and OTF-AL-fracture datasets, we employ ensemble learning to calculate the standard deviations of atomic forces for critical snapshots during OTF-AL of fracture simulations across four crack systems. These calculations are performed using five independent MTPs trained on datasets B+P+OTF-AL-GB and B+P+OTF-AL-GB+EIP-GS-fracture+OTF-AL-fracture, respectively. The results, presented in Figs S2i-p, reveal significant uncertainty at crack tip surfaces and dislocations nucleated from these tips before including crack-related LAEs. However, after incorporating the EIP-GS-fracture and OTF-AL-fracture datasets, the uncertainty for all atoms is substantially reduced, as evidenced by lower standard deviations in atomic forces. This improvement highlights the effectiveness of the combined EIP-GS and PCC-GCMC approaches in adapting to new applications, such as fracture simulations. These findings demonstrate that the final datasets, enriched with a comprehensive defect genome, provide robust capabilities for simulating general plastic deformations and crack propagation in BCC W, offering a solid foundation for accurate modeling across diverse loading configurations and boundary conditions.
Supplementary Note 4. Final MLIP training and selection
@@
We possess a comprehensive final database that aggregates representative LAEs in periodic configurations: B+P+OTF-AL-GB+EIP-GS-fracture+OTF-AL-fracture. Detailed breakdowns of each dataset are presented in Table S1. Altogether, the database encompasses 1026 configurations and 94,516 atoms. Of the 348 configurations newly incorporated, 276 are obtained through EIP-GS, while only 72 results from on-the-fly AL. This distribution highlights the efficiency of EIP-GS in significantly reducing the reliance on the more resource-intensive OTF-AL, thereby expediting the development of MLIPs.
@@
The MLIPs trained from this database are equipped to simulate nearly all plastic deformations and fracture of BCC W. However, configurations necessary for radiation simulations due to cascade displacement are currently not included. These configurations could easily be integrated from existing databases or generated through EIP-GS and OTF-AL in future enhancements. We utilize different levels of MTP and a complex ACE potential for training various MLIPs. We particularly favor ACE due to its computational efficiency and the ease of calculating extrapolation grades during simulations in LAMMPS, which greatly facilitates the ongoing assessment. The analysis of basic properties of BCC W, computational efficiency, and training/test performance are detailed in Table S2. It is observed that the accuracy of MTP models increases with their complexity level, but at the cost of reduced computational efficiency. Specifically, the computational expense of MTP20 is seven times that of MTP12. In contrast, ACE demonstrates accuracy comparable to MTP20 while maintaining efficiency similar to MTP12. Based on these findings, ACE is selected as the optimal model for modeling general defects in BCC W due to its superior balance of accuracy and efficiency.
@@
$N_{\text{cfg}}$ Average atom number Total atom number
Basic 678 72 48816
Initial EIP-GS 262 133 34846
OTF-AL-GB 47 135 6345
EIP-GS-fracture 14 118 1652
OTF-AL-fracture 25 115 2875
Total 1026 92 94392
@@
Figure S2: EIP-GS and OTF-AL of fracture simulations. a-h Simulation results of four crack systems using EAM and MEAM potentials. i-l Uncertainty quantification of defects near the crack tip based on MLIPs trained with B+P+OTF-AL-GB. m-p Uncertainty quantification of defects near the crack tip based on MLIPs trained with B+P+OTF-AL-GB+EIP-GS-fracture+OTF-AL-fracture. All atoms in a-h are color-coded by common neighbor analysis: blue atoms have a BCC lattice, green atoms have an FCC lattice, and white atoms have an unrecognized lattice. All atoms in i-p are color-coded by the standard deviation of atomic force.